A company will pay a dividend of  $1.25 next year. The dividend is expected to grow at a constant rate of 4.5% into the foreseeable future. The current stock price is $22.90 per share.  What are the stock’s expected dividend yield AND its expected total return for the coming year?

a.

5.46% and 9.96%

b.

4.26% and 8.76%

c.

6.25% and 10.75%

d.

4.26% and 4.50%

e.

5.46% and 4.50%
Answer #1

The Gordon growth formula is given as :

Price = Div/(R - g)

22.9 = 1.25/(R - 0.045)

Hence, R = 9.958%.

Expected Dividend yield will be = 1.25/22.9 = 5.458%

Hence, option A is correct.